Determine the equation of the graph of the given pa 2 axis of symmetry + 2 4 6 focus parabola -2+ vertex directrix -4
Answer #1

go ands of symmetry: pe-320 3 upward opening parabola -3 -240 phrass focus = (3-1) Vertex = (3,-2) -37 eqof tangent at the ve